### v2

@@ 2.2.2 Riemannian geometry of $\mathcal{L}_{p}^{++}$ and $\mathbb{P}(\mathcal{L}_{p}^{++})$
We review the Riemannian geometry of $\mathcal{L}_{p}^{++}$ and $\mathbb{P}(\mathcal{L}_{p}^{++})$ under Choleksy metric $g^{\text{Chol}}$ [lin2019]. The tangent spaces of each manifold are given by
@@
Proof 2.4.
See Appendix A.1 for the proof.
@@
Proof 2.5.
See Appendix A.1 for the proof.
It then directly follows that the Riemannian gradient and Hessian operator of a smooth function $f$ on $\mathbb{P}(\mathcal{L}_{p}^{++})$ are obtained as the orthogonal projections of those on $\mathcal{L}_{p}^{++}$ given in Proposition 2.
@@ 2.3 Quotient manifold
In this section, we review the quotient geometry of a Riemannian manifold. We again refer to [lee2012, lee2018] for the details. Suppose $(\mathcal{M},g)$ is a Riemannian manifold and $G$ is a Lie group acting smoothly, properly, and freely on $\mathcal{M}$. The action $(g,x)\in G\times\mathcal{M}\rightarrow g\cdot x\in\mathcal{M}$ is smooth and proper if it is smooth and proper as a map. Note that the map $f:X\rightarrow Y$ between two topological spaces is proper if the preimage of every compact subset of $Y$ is also compact in $X$. Also, the action is free if there is no non-trivial action that fixes the elements of $\mathcal{M}$, i.e., if $g\cdot x=x$, then $g$ is an identity $e$ for any $x\in\mathcal{M}$. If the Lie group $G$ is also compact, e.g., $\mathcal{O}_{p}$, then every smooth action of $G$ is proper ([lee2012], Corollary $21.6$). For any Lie group $G$ with a smooth, proper, and free action, there exists a unique smooth structure on $\mathcal{M}^{0}=\mathcal{M}/G$ such that the canonical projection $\pi:x\in\mathcal{M}\rightarrow[x]\in\mathcal{M}^{0}$ is a smooth submersion. Also, $\dim\mathcal{M}^{0}=\dim\mathcal{M}-\dim G$. With a submanifold $\mathcal{M}_{x}:=\pi^{-1}([x])$ of $\mathcal{M}$, the vertical space at $x$ is given as
@@
2.4 Review of algebraic geometry
In this section, we briefly review the algebraic geometry, focusing on the ingredients necessary for proving that the set of canonically decomposable matrices is Zariski-closed and has a Lebesgue measure zero in Section 4.1. We shall refer to [mumford1988, harris1992, hartshorne1977] for a more comprehensive review. For the subset $X\subset\real^{n}$, we say $X$ is Zariski-closed if $X$ is a zero locus of finitely many polynomials over the field . That is, for finitely many polynomials $p_{1},\ldots,p_{m}$,
@@ 2.4 Review of algebraic geometry
Otherwise, $X$ is Zariski-open. Note that such a set $X$ is also referred to as an affine (resp. projective) algebraic set with the affine space n (resp. the projective space $\real\mathbb{P}^{n-1}$). For a projective algebraic set, the polynomials should be homogeneous, i.e., each term has the same degree. To define the (topological) dimension of any subset $X$ of n, suppose $Y$ is a closed subset of n. We say $Y$ is reducible if $Y$ is the union of two proper closed subsets $Y_{1}$ and $Y_{2}$. Otherwise, $Y$ is irreducible. Then the (topological) dimension of $X$ is defined to be the largest integer $d\in[n]$ such that there exists a chain $Y_{0}\subsetneq Y_{1}\subsetneq\cdots\subsetneq Y_{d}\subset\bar{X}$, where each $Y_{i}$ is an irreducible closed subset of $\bar{X}$, the closure of $X$. Such $d$ always exists, and write $d:=\dim X$. We provide some useful facts about the topological dimension and Zariski-closed set to prove that the set of canonically decomposable matrices is proper Zariski-closed subset with measure zero, whose proofs are omitted (see Lemma 2.22.22.32.3 and $2.7$ of [domanov2015]).
@@
Lastly, we define the affine variety, the projective variety, and the Grassmannian over the field . The affine (resp. projective) variety is an irreducible affine (resp. projective) algebraic subset. Note that a product of affine (resp. projective) varieties $V_{1}$ and $V_{2}$ is again an affine (resp. projective) variety [sutherland2013]. The projective variety is known to be a complete variety; namely, the following is true for the projective variety [sutherland2013, mumford1988, harris1992].
@@
The Grassmannian $\text{Gr}(d,n)$ is a collection of $d$-dimensional linear subspaces of n, which is a manifold of dimension d(nd)d(n-d). While $\text{Gr}(d,n)$ can be realized in terms of both affine and projective varieties [devriendt2024], we focus on its projective variety aspect. Note that every $d$-dimensional linear subspace $D$ of n can be represented as a $d\times n$ matrix $Z$ whose rows represent the basis of $D$. Via the Plüker embedding [michalek2021], which realizes $\text{Gr}(d,n)$ as a projective variety, the matrix $Z$ can be identified with Plüker coordinates and the system of the polynomials that these coordinates should satisfy, so-called the Plüker equation [michalek2021] (see Example $1.1$ of [devriendt2024] for instance). Furthermore, if $\mathcal{P}_{Z}$ is an orthogonal projection onto $R(Z)$, then each entry of $|ZZ^{\top}|\mathcal{P}_{Z}$ is a quadratic polynomial in these Plüker coordinates ([devriendt2024], Theorem $2.1$).
@@ 3 Smooth manifold $\mathcal{C}_{p_{1},p_{2}}^{++}$
(6) $\begin{split}\mathcal{H}_{p_{1},p_{2},r}&:=\left\{A\in(\real^{p_{1}\times p_{2}})_{*}^{r}:\operatorname{\text{rank}}(A_{R})=p_{1},\operatorname{\text{rank}}(A_{C})=p_{2}\right\},\\ \mathcal{D}_{p_{1},p_{2},r}&:=F_{p_{1},p_{2},r}^{-1}(\left\{(I_{p_{1}}/p_{1},I_{p_{2}}/p_{2},p)\right\}),\;\mathcal{C}_{p_{1},p_{2},r}:=\varphi_{p_{1},p_{2},r}(\mathcal{D}_{p_{1},p_{2},r}),\\ F_{p_{1},p_{2},r}&:A\in\mathcal{H}_{p_{1},p_{2},r}\rightarrow(A_{R}/\textsf{tr}(A_{R}),A_{C}/\textsf{tr}(A_{R}),\textsf{tr}(A_{R}))\in\mathcal{R}_{p_{1},p_{2}},\\ \mathcal{R}_{p_{1},p_{2}}&:=\bar{\mathcal{S}}_{p_{1}}^{++}\times\bar{\mathcal{S}}_{p_{2}}^{++}\times\real_{+},\\ s_{p_{1},p_{2},r}&:[B]\in\real_{*}^{p\times r}/\mathcal{O}_{r}\rightarrow BB^{\top}\in\mathcal{S}_{p,r}^{+},\end{split}$
@@
We outline the proof strategy as follows. Although we state the strategy when $r=p$, note that this strategy can be straightforwardly extended to the rank-deficient case. Define the action of $\mathcal{O}_{p}$ on $\real{}_{*}^{p\times p}$ by $(O,B)\in\mathcal{O}_{p}\times\real_{*}^{p\times p}\rightarrow BO\in\real_{*}^{p\times p}$. Since the action is smooth, free, and proper as $\mathcal{O}_{p}$ is a Lie compact group, $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$ is a quotient manifold. To show that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a smooth manifold, observe that if $\mathcal{D}_{p_{1},p_{2}}$ is a smooth submanifold embedded in the smooth manifold $\mathcal{H}_{p_{1},p_{2}}$, so is $\mathcal{C}_{p_{1},p_{2}}$ in $\real{}_{*}^{p\times p}$ as the map $\varphi_{p_{1},p_{2}}$ is a diffeomorphism. Also, if $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant with the action above, we can show that $\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p_{1},p_{2}}$ is embeddded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$. The result that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is embedded in $\mathcal{S}_{p_{1},p_{2}}^{++}$ then follows from the facts that the map $s_{p_{1},p_{2}}$ is a diffeomorphism ([massart2020], Proposition $2.8$) and $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv s_{p_{1},p_{2}}(\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p})$. This strategy and the ancillary results below can be applied when $r<p$. The only difference is the way to show $\mathcal{D}_{p_{1},p_{2},r}$ is a smooth manifold as shown in Section 4.2. Now taking $r=p$, we provide the ancillary results to prove the main result of this section.
@@
Proof 3.2.
Let $\phi_{R}:A\in(\real^{p_{1}\times p_{2}})_{*}^{p}\rightarrow[A_{1},\ldots,A_{p}]\in\real^{p_{1}\times pp_{2}}$ and $\phi_{C}:A\in(\real^{p_{1}\times p_{2}})_{*}^{p}\rightarrow[A_{1}^{\top},\ldots,A_{p}^{\top}]\in\real^{p_{2}\times pp_{1}}$. For any $A\in(\real^{p_{1}\times p_{2}})_{*}^{p}$, $A_{R}\equiv\phi_{R}(A)\phi_{R}(A)^{\top}$ and $A_{C}\equiv\phi_{C}(A)\phi_{C}(A)^{\top}$ are of full-rank if and only if $\phi_{R}(A)$ and $\phi_{C}(A)$ are so, respectively. Thus, $\mathcal{H}_{p_{1},p_{2}}=\phi_{R}^{-1}(\real_{*}^{p_{1}\times pp_{2}})\cap\phi_{C}^{-1}(\real_{*}^{p_{2}\times pp_{1}})$. Note that both the maps $\phi_{R}$ and $\phi_{C}$ are smooth, and $\real{}_{*}^{p_{1}\times pp_{2}}$ and $\real{}_{*}^{p_{2}\times pp_{1}}$ are open in their respective ambient space. Thus, both $\phi_{R}^{-1}(\real_{*}^{p_{1}\times pp_{2}})$ and $\phi_{C}^{-1}(\real_{*}^{p_{2}\times pp_{1}})$ are open in $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Hence, the transversality theorem ([lee2012], Theorem 6.35) implies that $\mathcal{H}_{p_{1},p_{2}}$ is an open submanifold of $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Because $(\real^{p_{1}\times p_{2}})_{*}^{p}$ is diffeomorphic to $\real{}_{*}^{p\times p}$ via the map $\varphi_{p_{1},p_{2}}$, $(\real^{p_{1}\times p_{2}})_{*}^{p}$ is open in $(\real^{p_{1}\times p_{2}})^{p}$. Therefore, $T_{A}\mathcal{H}_{p_{1},p_{2}}=T_{A}(\real^{p_{1}\times p_{2}})_{*}^{p}\equiv(\real^{p_{1}\times p_{2}})^{p}$ for any $A\in\mathcal{H}_{p_{1},p_{2}}$.
@@
We provide a complete proof of Proposition 4 in Appendix A.2. Here we provide the main idea of the proof.
Proof 3.3 (Sketch of Proof).
We use the constant-rank level set theorem ([lee2012], Theorem $5.12$) to prove the result. Take $B=(B_{1},\ldots,B_{p})\in T_{A}\mathcal{H}_{p_{1},p_{2}}$. Let $a=[\text{vec}(A_{1})^{\top},\ldots,\text{vec}(A_{p})^{\top}]^{\top}$, and $b=[\text{vec}(B_{1})^{\top},\ldots,\text{vec}(B_{p})^{\top}]^{\top}$. Since $\textsf{tr}(A_{R})=p$ for any $A\in\mathcal{D}_{p_{1},p_{2}}$, the differential of $F_{p_{1},p_{2}}$ at $A$ is given by
@@
(13) $\left[\begin{array}[]{c}J_{1}\\ J_{2}\\ J_{3}\end{array}\right]b:=\underbrace{\left[\begin{array}[]{c}\frac{1}{p}(I_{p_{1}^{2}}+K_{(p_{1},p_{1})})[A_{1}\otimes I_{p_{1}},\ldots,A_{p}\otimes I_{p_{1}}]-\frac{2}{p_{1}^{2}p_{2}}\text{vec}(I_{p_{1}})a^{\top}\\ \frac{1}{p}(I_{p_{2}^{2}}+K_{(p_{2},p_{2})})[I_{p_{2}}\otimes A_{1}^{\top},\ldots,I_{p_{2}}\otimes A_{p}^{\top}]-\frac{2}{p_{1}p_{2}^{2}}\text{vec}(I_{p_{2}})a^{\top}\\ 2a^{\top}\end{array}\right]}_{J:=J(A)}b.$
@@
Then we claim in Appendix A.2 that
(14) $\begin{split}&\dim C(J_{1}^{\top})=\binom{p_{1}+1}{2}-1,\,\dim C(J_{2}^{\top})=\binom{p_{2}+1}{2}-1,\,\dim C(J_{3}^{\top})=1,\\ &\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})=\dim C(J_{2}^{\top})\cap C(J_{3}^{\top})=\dim C(J_{1}^{\top})\cap C(J_{3}^{\top})\\ =&\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})\cap C(J_{3}^{\top})=0.\end{split}$
@@
By Proposition 4, the image of $\mathcal{D}_{p_{1},p_{2}}$ by the diffeomorphism $\varphi_{p_{1},p_{2}}$, $\mathcal{C}_{p_{1},p_{2}}$, is closed and embedded in $\real{}_{*}^{p\times p}$. As discussed above, we show that the smooth manifold $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant, and closed and embedded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$.
@@
Proof 3.5.
where the last equality holds as $(X_{1},\ldots,X_{p})\in\mathcal{D}_{p_{1},p_{2}}$. Similarly, $\tilde{Y}_{C}\tilde{Y}_{C}^{\top}=p_{1}I_{p_{2}}$.
@@
Proof 3.7.
By the quotient manifold theorem ([lee2012], Theorem $21.10$), both $\mathcal{M}/G$ and $\mathcal{N}/G$ are smooth manifolds. Let $i:\mathcal{N}/G\xhookrightarrow{}\mathcal{M}/G$ be an inclusion map. We claim that $i$ is a smooth embedding. By Theorem $4.4$ of [martin2017], $i$ is an injective immersion. To show that $i$ is a topological embedding, let $\pi:X\in\mathcal{M}\rightarrow[X]\in\mathcal{M}/G$ be the canonical projection, which is a smooth submersion. Hence, $\pi$ is open and $\mathcal{N}$ is $G$-invariant so that $\mathcal{N}$ is saturated for the map $\pi$. Here the subset $C\subset U$ is saturated for the map $f:U\rightarrow V$ between two topological spaces if $f^{-1}(f(C))=C$. Since $\mathcal{N}$ is a topological subspace of $\mathcal{M}$ as it is embedded, so is $\pi(\mathcal{N})\equiv\mathcal{N}/G$ of the quotient space $\pi(\mathcal{M})\equiv\mathcal{M}/G$ ([munkres2000], Theorem $22.1$). Thus, $i$ is a topological embedding.
@@
Proof 3.8.
The compactness follows as (3) implies that $\textsf{tr}(C)=p$ for any $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$. To show that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a smooth submanifold embedded in $\mathcal{S}_{p}^{++}$, note that $\mathcal{D}_{p_{1},p_{2}}$ is embedded in $\mathcal{H}_{p_{1},p_{2}}$ by Proposition 4, and $\mathcal{H}_{p_{1},p_{2}}$ is also embedded in $(\real^{p_{1}\times p_{2}})_{*}^{p}$ as an open submanifold. Thus, $\mathcal{D}_{p_{1},p_{2}}$ is embedded in $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Hence, $\mathcal{C}_{p_{1},p_{2}}\equiv\varphi_{p_{1},p_{2}}(\mathcal{D}_{p_{1},p_{2}})$ is embedded in $\real{}_{*}^{p\times p}\equiv\varphi_{p_{1},p_{2}}((\real^{p_{1}\times p_{2}})_{*}^{p})$ as the map $\varphi_{p_{1},p_{2}}$ is a diffeomorphism. By Lemma 3.4, $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant. Thus, taking $\mathcal{M}=\real_{*}^{p\times p}$, $\mathcal{N}=\mathcal{C}_{p_{1},p_{2}}$, and $G=\mathcal{O}_{p}$ in Lemma 3.6, we have that $\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p}$ is embedded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$. Also, the quotient manifold theorem implies that
@@
By Lemma 3.4 and (3), we have that $AA^{\top}=BB^{\top}\in\mathcal{C}_{p_{1},p_{2}}^{++}$ for $A,B\in\real_{*}^{p\times p}$ if and only if $A,B\in\mathcal{C}_{p_{1},p_{2}}$ and $A=BO$ for some $O\in\mathcal{O}_{p}$. Because the map $s_{p_{1},p_{2}}$ defined in (6) is a diffeomorphism by Proposition $2.8$ of [massart2020], $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv s_{p_{1},p_{2}}(\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p})$ is a smooth submanifold embedded in $\mathcal{S}_{p}^{++}\equiv s_{p_{1},p_{2}}(\real_{*}^{p\times p}/\mathcal{O}_{p})$.
@@
Proof 3.9.
(15) $\textsf{tr}_{1}(W)=\mathbf{0}_{p_{1}\times p_{1}},\quad\textsf{tr}_{2}(W)=\mathbf{0}_{p_{2}\times p_{2}}.$

### v3

@@ 2.2.2 Riemannian geometry of $\mathcal{L}_{p}^{++}$ and $\mathbb{P}(\mathcal{L}_{p}^{++})$
We review the Riemannian geometry of $\mathcal{L}_{p}^{++}$ and $\mathbb{P}(\mathcal{L}_{p}^{++})$ under Choleksy metric $g^{\text{Chol}}$ [40]. The tangent spaces of each manifold are given by
@@
Proof.
See Appendix A.1 for the proof. ∎
@@
Proof.
See Appendix A.1 for the proof. ∎
It then directly follows that the Riemannian gradient and Hessian operator of a smooth function $f$ on $\mathbb{P}(\mathcal{L}_{p}^{++})$ are obtained as the orthogonal projections of those on $\mathcal{L}_{p}^{++}$ given in Proposition 2.3.
@@ 2.3 Quotient manifold
In this section, we review the quotient geometry of a Riemannian manifold. We again refer to [38, 39] for the details. Suppose $(\mathcal{M},g)$ is a Riemannian manifold and $G$ is a Lie group acting smoothly, properly, and freely on $\mathcal{M}$. The action $(g,x)\in G\times\mathcal{M}\rightarrow g\cdot x\in\mathcal{M}$ is smooth and proper if it is smooth and proper as a map. Note that the map $f:X\rightarrow Y$ between two topological spaces is proper if the preimage of every compact subset of $Y$ is also compact in $X$. Also, the action is free if there is no non-trivial action that fixes the elements of $\mathcal{M}$, i.e., if $g\cdot x=x$, then $g$ is an identity $e$ for any $x\in\mathcal{M}$. If the Lie group $G$ is also compact, e.g., $\mathcal{O}_{p}$, then every smooth action of $G$ is proper ([38], Corollary $21.6$). For any Lie group $G$ with a smooth, proper, and free action, there exists a unique smooth structure on $\mathcal{M}^{0}=\mathcal{M}/G$ such that the canonical projection $\pi:x\in\mathcal{M}\rightarrow[x]\in\mathcal{M}^{0}$ is a smooth submersion. Also, $\dim\mathcal{M}^{0}=\dim\mathcal{M}-\dim G$. With a submanifold $\mathcal{M}_{x}:=\pi^{-1}([x])$ of $\mathcal{M}$, the vertical space at $x$ is given as
@@
2.4 Algebraic geometry
In this section, we briefly review the algebraic geometry, focusing on the ingredients necessary for proving that the set of canonically decomposable matrices is Zariski-closed and has a Lebesgue measure zero in Section 4.1. We shall refer to [49, 31, 32] for a more comprehensive review. For the subset $X\subset\real^{n}$, we say $X$ is Zariski-closed if $X$ is a zero locus of finitely many polynomials over the field . That is, for finitely many polynomials $p_{1},\ldots,p_{m}$,
@@ 2.4 Algebraic geometry
[…]
@@
Lastly, we define the affine variety, the projective variety, and the Grassmannian over the field . The affine (resp. projective) variety is an irreducible affine (resp. projective) algebraic subset. Note that a product of affine (resp. projective) varieties $V_{1}$ and $V_{2}$ is again an affine (resp. projective) variety [64]. The projective variety is known to be a complete variety; namely, the following is true for the projective variety [64, 49, 31].
@@
The Grassmannian $\text{Gr}(d,n)$ is a collection of $d$-dimensional linear subspaces of n, which is a manifold of dimension d(nd)d(n-d). While $\text{Gr}(d,n)$ can be realized as both affine and projective varieties [20], we focus on its projective variety aspect. Note that every $d$-dimensional linear subspace $D$ of n can be represented as a $d\times n$ matrix $Z$ whose rows represent the basis of $D$. Via the Plüker embedding [47], which realizes $\text{Gr}(d,n)$ as a projective variety, the matrix $Z$ can be identified with Plüker coordinates and the system of the polynomials that these coordinates should satisfy, so-called the Plüker equation [47] (see Example $1.1$ of [20] for instance). Furthermore, if $\mathcal{P}_{Z}$ is an orthogonal projection onto $R(Z)$, then each entry of $|ZZ^{\top}|\mathcal{P}_{Z}$ is a quadratic polynomial in these Plüker coordinates ([20], Theorem $2.1$).
@@ 3 Smooth manifold $\mathcal{C}_{p_{1},p_{2}}^{++}$
$\begin{split}\mathcal{H}_{p_{1},p_{2},r}&:=\left\{A\in(\real^{p_{1}\times p_{2}})_{*}^{r}:\operatorname{\text{rank}}(A_{R})=p_{1},\operatorname{\text{rank}}(A_{C})=p_{2}\right\},\\ \mathcal{D}_{p_{1},p_{2},r}&:=F_{p_{1},p_{2},r}^{-1}(\left\{(I_{p_{1}}/p_{1},I_{p_{2}}/p_{2},p)\right\}),\;\mathcal{C}_{p_{1},p_{2},r}:=\varphi_{p_{1},p_{2},r}(\mathcal{D}_{p_{1},p_{2},r}),\\ F_{p_{1},p_{2},r}&:A\in\mathcal{H}_{p_{1},p_{2},r}\rightarrow(A_{R}/\textsf{tr}(A_{R}),A_{C}/\textsf{tr}(A_{R}),\textsf{tr}(A_{R}))\in\mathcal{R}_{p_{1},p_{2}},\\ \mathcal{R}_{p_{1},p_{2}}&:=\bar{\mathcal{S}}_{p_{1}}^{++}\times\bar{\mathcal{S}}_{p_{2}}^{++}\times\real_{+},\\ s_{p_{1},p_{2},r}&:[B]\in\real_{*}^{p\times r}/\mathcal{O}_{r}\rightarrow BB^{\top}\in\mathcal{S}_{p,r}^{+},\end{split}$ (7)
@@
We outline the proof strategy as follows. Although we state the strategy when $r=p$, note that this strategy can be straightforwardly extended to the rank-deficient case. Define the action of $\mathcal{O}_{p}$ on $\real{}_{*}^{p\times p}$ by $(O,B)\in\mathcal{O}_{p}\times\real_{*}^{p\times p}\rightarrow BO\in\real_{*}^{p\times p}$. Since the action is smooth, free, and proper as $\mathcal{O}_{p}$ is a Lie compact group, $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$ is a quotient manifold. To show that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a smooth manifold, observe that if $\mathcal{D}_{p_{1},p_{2}}$ is a smooth submanifold embedded in the smooth manifold $\mathcal{H}_{p_{1},p_{2}}$, so is $\mathcal{C}_{p_{1},p_{2}}$ in $\real{}_{*}^{p\times p}$ as the map $\varphi_{p_{1},p_{2}}$ is a diffeomorphism. Also, if $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant with the action above, we can show that $\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p_{1},p_{2}}$ is embeddded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$. The result that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is embedded in $\mathcal{S}_{p_{1},p_{2}}^{++}$ then follows from the facts that the map $s_{p_{1},p_{2}}$ is a diffeomorphism ([45], Proposition $2.8$) and $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv s_{p_{1},p_{2}}(\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p})$. This strategy and the ancillary results below can be applied when $r<p$. The only difference is the way to show $\mathcal{D}_{p_{1},p_{2},r}$ is a smooth manifold as shown in Section 4.2. Now taking $r=p$, we provide the ancillary results to prove the main result of this section.
@@
Proof.
See Appendix A.2 for the proof. ∎
@@
We provide a complete proof of Proposition 3.2 in Appendix A.2. Here we provide the main idea of the proof.
Sketch of Proof.
We use the constant-rank level set theorem ([38], Theorem $5.12$) to prove the result. Take $B=(B_{1},\ldots,B_{p})\in T_{A}\mathcal{H}_{p_{1},p_{2}}$. Let $a=[\text{vec}(A_{1})^{\top},\ldots,\text{vec}(A_{p})^{\top}]^{\top}$, and $b=[\text{vec}(B_{1})^{\top},\ldots,\text{vec}(B_{p})^{\top}]^{\top}$. Since $\textsf{tr}(A_{R})=p$ for any $A\in\mathcal{D}_{p_{1},p_{2}}$, the differential of $F_{p_{1},p_{2}}$ at $A$ is given by
@@
$\left[\begin{array}[]{c}J_{1}\\ J_{2}\\ J_{3}\end{array}\right]b:=\underbrace{\left[\begin{array}[]{c}\frac{1}{p}(I_{p_{1}^{2}}+K_{(p_{1},p_{1})})[A_{1}\otimes I_{p_{1}},\ldots,A_{p}\otimes I_{p_{1}}]-\frac{2}{p_{1}^{2}p_{2}}\text{vec}(I_{p_{1}})a^{\top}\\ \frac{1}{p}(I_{p_{2}^{2}}+K_{(p_{2},p_{2})})[I_{p_{2}}\otimes A_{1}^{\top},\ldots,I_{p_{2}}\otimes A_{p}^{\top}]-\frac{2}{p_{1}p_{2}^{2}}\text{vec}(I_{p_{2}})a^{\top}\\ 2a^{\top}\end{array}\right]}_{J:=J(A)}b.$ (14)
@@
We claim in Appendix A.2 that
$\begin{split}&\dim C(J_{1}^{\top})=\binom{p_{1}+1}{2}-1,\,\dim C(J_{2}^{\top})=\binom{p_{2}+1}{2}-1,\,\dim C(J_{3}^{\top})=1,\\ &\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})=\dim C(J_{2}^{\top})\cap C(J_{3}^{\top})=\dim C(J_{1}^{\top})\cap C(J_{3}^{\top})\\ =&\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})\cap C(J_{3}^{\top})=0.\end{split}$ (15)
@@
By Proposition 3.2, the image of $\mathcal{D}_{p_{1},p_{2}}$ by the diffeomorphism $\varphi_{p_{1},p_{2}}$, $\mathcal{C}_{p_{1},p_{2}}$, is closed and embedded in $\real{}_{*}^{p\times p}$. As discussed above, we show that the smooth manifold $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant, and closed and embedded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$.
@@
Proof.
See Appendix A.2 for the proof. ∎
@@
Proof.
See Appendix A.2 for the proof. ∎
@@
Proof.
The compactness follows as (4) implies that $\textsf{tr}(C)=p$ for any $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$. To show that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a smooth submanifold embedded in $\mathcal{S}_{p}^{++}$, note that $\mathcal{D}_{p_{1},p_{2}}$ is embedded in $\mathcal{H}_{p_{1},p_{2}}$ by Proposition 3.2, and $\mathcal{H}_{p_{1},p_{2}}$ is also embedded in $(\real^{p_{1}\times p_{2}})_{*}^{p}$ as an open submanifold. Thus, $\mathcal{D}_{p_{1},p_{2}}$ is embedded in $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Hence, $\mathcal{C}_{p_{1},p_{2}}\equiv\varphi_{p_{1},p_{2}}(\mathcal{D}_{p_{1},p_{2}})$ is embedded in $\real{}_{*}^{p\times p}\equiv\varphi_{p_{1},p_{2}}((\real^{p_{1}\times p_{2}})_{*}^{p})$ as the map $\varphi_{p_{1},p_{2}}$ is a diffeomorphism. By Lemma 3.3, $\mathcal{C}_{p_{1},p_{2}}$ is $\mathcal{O}_{p}$-invariant. Thus, taking $\mathcal{M}=\real_{*}^{p\times p}$, $\mathcal{N}=\mathcal{C}_{p_{1},p_{2}}$, and $G=\mathcal{O}_{p}$ in Lemma 3.4, we have that $\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p}$ is embedded in $\real{}_{*}^{p\times p}/\mathcal{O}_{p}$. Also, the quotient manifold theorem implies that
@@
By Lemma 3.3 and (4), we have that $AA^{\top}=BB^{\top}\in\mathcal{C}_{p_{1},p_{2}}^{++}$ for $A,B\in\real_{*}^{p\times p}$ if and only if $A,B\in\mathcal{C}_{p_{1},p_{2}}$ and $A=BO$ for some $O\in\mathcal{O}_{p}$. Because the map $s_{p_{1},p_{2}}$ defined in (7) is a diffeomorphism by Proposition $2.8$ of [45], $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv s_{p_{1},p_{2}}(\mathcal{C}_{p_{1},p_{2}}/\mathcal{O}_{p})$ is a smooth submanifold embedded in $\mathcal{S}_{p}^{++}\equiv s_{p_{1},p_{2}}(\real_{*}^{p\times p}/\mathcal{O}_{p})$. ∎
@@
Proof.
See Appendix A.2 for the proof. ∎